5n - 16 - 8n = -105n - 16 - 8n = -10 please help me
Answer by stanbon(75887)

5n - 16 - 8n = -105n - 16 - 8n = -10
-------------
Comment::
You are mixing two equations.
Do you want 5n-16 = -10
or do you want -105n-16-8n = -10 ????
Cheers,
Stan H.
-------------